# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt

### v1

@@ 5.4 Ablation Study
Impact of Constraints during Optimization. In the absence of ground truth labels, it is crucial to set appropriate targets for testing-time optimization to achieve high-quality 3D poses. To this end, we evaluate the effectiveness of two proposed constraints in the optimization process. Table 3 presents the ablation results of 5 iterations using the UAO framework on the test set of Human3.6M. The results reveal that the projection constraint leads to a significant performance improvement of the pure data-driven method (from 49.5mm to 48.2mm). Furthermore, the introduction of the uncertainty constraint brings further improvements in MPJPE, with the metric reaching 47.2mm. These experimental results demonstrate the effectiveness of these two constraints. To further explore the optimization mechanism, we illustrate the performance with different iterations on the test set of Human3.6M, as shown in Figure 4. We observe that using only the projection constraint leads to a rapid decrease in MPJPE, followed by a gradual increase. In contrast, the error can be stabilized at a lower level with the further introduction of uncertainty constraint, even with more iterations. The main reason is that the projection constraint only ensures the consistency of the projected pose in the 2D plane, which is easy to overfitting, yielding unrealistic poses in the 3D space. For complementary, the uncertainty constraint ensures the output remains close to the 3D manifold, which restricts joints with low uncertainty from deviating too far from the ground truth while allowing joints with high uncertainty to have more optimization flexibility. The experimental results showcase the robust adaptability of our method, especially in real-world scenarios where the optimal number of iterations is uncertain.
@@
Figure 5: Combination of a histogram representing the uncertainty for each joint and a line chart depicting the mean error (MPJPE) for each joint. The left y-axis denotes the magnitude of uncertainty, while the right y-axis signifies the error value.
Table 4: Inference speed of various optimization methods. UAO* signifies employing the network’s latent state as the variable for optimization.
Method FPS MPJPE $\downarrow$
Model-only 310 49.5
BOA [8] 19 48.5
ISO [41] 18 47.9
UAO 48 47.2
UAO* 125 47.7
Impact of Optimization Strategies. We evaluated the inference speed of various optimization strategies on a single frame using one GeForce RTX 3090, results are shown in Table 4. We use GUMLP as the baseline model, achieving 310 fps. By setting the number of iterations as 4, our UGO framework reaches 48 fps, outperforming the speed of both the teacher-student approach used in BOA [8] and the Inference Stage Optimization (ISO) [41] approach. This is because our UAO framework optimize the latent state $\mathbf{z}$ rather than the entire model parameters, resulting in accelerated inference. Despite this reduction, the performance still meets real-time requirements. It is worth noting that the output of any layer within the network can be selected as the latent state for optimization. Here, we select the latent state before the final fully connected layer as the optimized variable, termed UAO*. This choice brings a significant surge in inference speed, reaching up to 125 fps. Such an alteration allows faster processing while maintaining satisfactory performance.
@@
Figure 6: Visualization of the iterative optimization process. The red pose is the ground truth and the blue pose is our prediction. Color intensity reflects the number of iterations, with darker colors indicating more iterations. The initial 3D pose generated by the pre-trained model (column 3) progressively improves through iterative optimization (columns 4-7).
5.5 Uncertainty and Error for each Joint
@@
5.6 Generalization to Unseen Camera Views
Table 5: Performance across different views on Human3.6M in MPJPE. In-domain refers to testing under the same view as the training, while out-of-domain refers to testing under views other than the training view. $P_{i}$ denotes the $i$th camera view.
@@ 5.6 Generalization to Unseen Camera Views
Table 5 presents the relative gains of the proposed approach over in- and out-domain cases on Human3.6M, where four camera views are available. It can be inferred that the UAO framework can not only help the in-domain cases but also bring more significant improvements to unseen camera views. This verifies its ability to break the domain gap and meet our design expectations. Besides, from the last column of the table, a large training set can bring more optimization gains, which may be explained by better pre-training leading to a better depiction of the 3D pose manifold.
5.7 Qualitative Results
@@
The visualization results on the Human3.6M dataset are shown in Figure 7. We can see that our original GUMLP is marginally superior to the previous state-of-the-art method, MGCN [47]. Moreover, employing the UAO framework brings better generation with more precise poses. For example, the 3D pose of photoing action (row 2) predicted by GUMLP with UAO (col 4) is similar to the ground truth. As a comparison, the left arms predicted by MGCN and the GUMLP are at a lower position. We further visualize the iterative process of the UAO framework, as shown in Figure 6. After four iterations of optimization, the original 3D output of GUMLP (col 3) becomes more reasonable and accurate in detail, highlighting the effectiveness of the optimization process. Figure 8 shows the qualitative results on MPI-INF-3DHP test set. It can be seen that our approach exhibits commendable performance in both indoor and outdoor scenarios.
Figure 8: Qualitative results of our method on indoor and outdoor scenes on the MPI-INF-3DHP dataset.
6 Conclusion
@@
In this paper, we propose an Uncertainty-Aware testing-time Optimization (UAO) framework for 3D human pose estimation. During the training process, we propose the GUMLP to estimate 3D results and uncertainty values for each joint. For test-time optimization, our UAO framework freezes the pre-trained network parameters and optimizes a latent state initialized by the input 2D pose. To constrain the optimization direction in both 2D and 3D spaces, projection and uncertainty constraints are applied. Extensive experiments show that our approach achieves state-of-the-art performance on two popular datasets.
Supplementary Material
This supplemental material contains the following parts:
  • Theory of uncertainty estimation (Sec. A).
  • Brief introduction of GCN (Sec. B).
  • Additional quantitative results (Sec. C).
  • Additional ablation studies (Sec. D).
  • Effectiveness of the UAO framework (Sec. E).
  • Limitations (Sec. F).
Appendix A Theory of Uncertainty Estimation
Assuming the 3D pose follows Gaussian distributions, we reformulate the 2D-to-3D network as a Bayesian Neural Network (BNN) [13], which estimates a Gaussian distribution of the target rather than only predicting the absolute coordinates of points. Specifically, for the 3D joints $\boldsymbol{J}_{i}$ of person $i$ in the dataset, our goal is to estimate the distribution of 3D joints for that individual. During training, the objective is to maximize the likelihood of all observation pairs. The likelihood function can be formulated as follows:
$Likelihood=\prod_{i,k}\mathcal{N}(\boldsymbol{J}_{i,k}|\boldsymbol{\mu}_{i,k},\Sigma_{i,k}),$ (7)
where $k$ is the index for the $k$-th joint of a person, $\boldsymbol{\mu}_{i,k}$ represents the mean, $\Sigma_{i,k}$ represents the covariance matrix of the Gaussian distribution for the $k$-th joint of person $i$. Each joint is considered as a random variable following a normal Gaussian distribution. To further elaborate, the mean $\boldsymbol{\mu}_{i,k}$ captures the most probable location of the $k$-th joint, while the covariance matrix $\Sigma_{i,k}$ characterizes the uncertainty of the distribution around this mean. The likelihood term models the probability of observing the actual 3D joint positions given the estimated means and covariances. In pursuit of the most accurate estimation of the 3D joint positions, we seek the maximum likelihood, which aligns the estimated Gaussian distribution with the observed 3D joint positions, capturing both the most probable locations $\boldsymbol{\mu}_{i,k}$ and the uncertainty $\Sigma_{i,k}$ associated with each joint. The formulation of the maximum likelihood function can be achieved by:
$\left.+\ln|\Sigma_{i,k}|\right).$ (8)
Here we simplify the covariance matrix $\Sigma_{i,k}$ as a diagonal matrix with diagonal elements $\boldsymbol{\sigma}_{i,k}^{2}$, then we get:
[…]
SimpleBaseline (ICCV’2017) [21] 39.5 43.2 46.4 47.0 51.0 56.0 41.4 40.6 56.5 69.4 49.2 45.0 49.5 38.0 43.1 47.7
VideoPose3D (CVPR’2019) [23] 36.0 38.7 38.0 41.7 40.1 45.9 37.1 35.4 46.8 53.4 41.4 36.9 43.1 30.3 34.8 40.0
LCN (ICCV’2019) [4] 36.9 41.6 38.0 41.0 41.9 51.1 38.2 37.6 49.1 62.1 43.1 39.9 43.5 32.2 37.0 42.2
[…]
SRNet (ECCV’2020) [37] 35.8 39.2 36.6 36.9 39.8 45.1 38.4 36.9 47.7 54.4 38.6 36.3 39.4 30.3 35.4 39.4
[…]
MLP-JCG (TMM’2023) [30] 33.7 37.4 37.3 39.6 39.8 47.1 33.7 33.8 45.7 60.5 39.7 37.7 40.1 30.1 33.8 39.3
GKONet (TCSVT’2023) [11] 35.4 38.8 35.9 40.4 39.6 44.0 36.7 35.4 46.8 53.7 40.9 36.6 42.0 30.6 33.9 39.4
Appendix B Graph Convolutional Network
Graph Convolutional Network (GCN) is capable to capture intricate relationships and structures within graph-structured data. Consider an undirected graph as $G{=}\{V,E\}$, where $V$ is the set of nodes, and $E$ is the set of edges. The edges can be encoded in an adjacency matrix $A{\in}\{0,1\}^{N{\times}N}$. For the input $X_{l}$ of the $l^{th}$ layer, the vanilla graph convolution aggregates the features of the neighboring nodes. The output $X_{l+1}$ of the $l^{th}$ GCN layer can be formulated as:
Table 7: Influence of projection loss ($\lambda_{P}$) and uncertainty loss ($\lambda_{U}$) hyperparameters.
$\lambda_{P}$ $\lambda_{U}$ Best Iteration MPJPE
1.0 0.000 3 47.6
0.005 5 47.2
0.010 8 47.5
0.015 10 47.9
0.020 9 48.4
Appendix C Quantitative Results under Protocol 2
Table 6 compares the performance of our approach with other state-of-the-art methods on Human3.6M under Protocol 2. Following previous approaches [47, 2, 35], the 2D pose detected by CPN [3] is used as input. The experimental results show that our GUMLP achieves competitive results. Furthermore, the performance of GUMLP under Protocol 2 can be further improved when using the UAO framework, surpassing all previous methods.
Appendix D Additional Ablation Studies
Impact of $\lambda_{P}$ and $\lambda_{U}$. During testing-time optimization, we applied constraints to ensure alignment of the generated poses in both 2D and 3D spaces. Table 7 shows the impact of the hyperparameters for the projection loss ($\lambda_{P}$) and uncertainty loss ($\lambda_{U}$). Here we set $\lambda_{P}=1$ for projection loss and explore the impact of the hyperparameter $\lambda_{U}$ for uncertainty loss. As the uncertainty loss regulates the degree to which the 3D pose updates can deviate from the initial estimation, a small or large value of $\lambda_{U}$ will influence the performance. For better performance, the final hyperparameters selected in our experiments are $\lambda_{P}=1.0$ and $\lambda_{U}=0.005$.
Figure 9: Performance with different iterations on Human3.6M test set. Here we adopt the pre-trained GUMLP as baseline and use the 2D pose detected by CPN [3] to initialize the latent state $\mathbf{z}$.
Appendix E Effectiveness of the UAO Framework
We have introduce the optimization strategy of the proposed UAO framework in Sec 4.1 of the main text. By freezing the pre-trained model parameters and selecting a hidden state $\mathbf{z}$ as the optimized variable, we establish a robust testing-time optimization approach. In this section, we further validate the effectiveness of UAO framework.
Algorithm 2 Online Adaptation of BOA
1:2D poses $\{\mathbf{J}_{j}^{2D}\}_{j=1}^{N}$ from test set, pre-trained model $M_{\phi_{0}}$ that parameterized by ${\phi_{0}}$, moving average rate $\alpha$ of the teacher model $M_{\omega}$ that parameterized by ${\omega}$, learning rate $\eta$, projection matrix $\mathbf{P}$, optimization steps $T$
2:3D pose $\hat{\mathbf{J}}_{j}^{3D}$
3:Initialization: ${\omega}={\phi_{0}}$
4:for $j=1$ to $N$ do
5:     $\phi_{j}=\phi_{j-1}$
6:     for $i$ = 0 to $T$ do
7:         $\hat{\mathbf{J}}_{j}^{3D}=M_{\phi_{j}}(\mathbf{J}_{j}^{2D})$
8:         $L^{opt}=\|\hat{\mathbf{J}}_{j}^{3D}\cdot\mathbf{P}-\mathbf{J}_{j}^{2D}\|_{2}$
9:         $\phi_{j}=\phi_{j}-\eta\cdot\nabla L^{opt}$
10:     end for
11:     # update the teacher model
12:     ${\omega}=(1-\alpha)\cdot{\omega}+\alpha\cdot\phi_{j}$
13:     # predict 3D pose by the teacher model
14:     $\hat{\mathbf{J}}_{j}^{3D}=M_{\omega}(\mathbf{J}_{j}^{2D})$
15:end for
Comparison with other optimization-based methods. We further compare our method with vanilla ISO [41] and BOA [8], two typical testing-time optimization methods. Both of them seek for better generalization capabilities through fine-tuning the network on the test set. The ISO adopts an early-stop mechanism, which requires reloading the pre-trained parameters each time new data is encountered. On the other hand, the BOA adopts a teacher model [31], requiring the loading of model parameters only once and continuously updating them through test data. The implementation details of BOA are summarized in Algorithm 2. Using the pre-trained GUMLP model as baseline, the comparison of the average performance across different iterations is illustrated in Figure 9. The results show that our UAO framework achieves the best performance, demonstrating its superiority. Additionally, the UAO framework optimizes only a hidden state $\mathbf{z}$ rather than the entire network parameters, thereby requiring the storage of a small number of parameters. As presented in Table 5 in the main text, UAO attains the highest inference speed, making it more suitable for online inference scenarios.
Extension to Other Data-Driven Models. To evaluate the general applicability of the UAO framework, we test its performance on other data-driven models. For simplicity, here we only adopt the projection constraint during testing-time optimization. The results are presented in Table 8. The outcomes clearly indicate a substantial performance enhancement for both multi-frame and single-frame models through the implementation of testing-time optimization paradigms. The experimental results underscore the generalization of our UAO framework, showcasing its adaptability and ease of extension to other pre-trained models.
Table 8: Performance of other data-driven models with UAO framework in MPJPE. MF represents multi-frame methods, and SF represents single-frame methods. GT denotes the method adopts ground truth 2D input.
Method Original + UAO $\Delta$
MF MHFormer [18] 43.0 40.8 2.2
MixSTE [42] 40.9 39.9 1.0
PoseFormer [46] 44.3 43.2 1.1
SF SimpleBaseline [21] 53.0 50.2 2.8
SemGCN [44] (GT) 42.6 42.0 0.6
MGCN [47] 49.4 47.9 1.5
Appendix F Limitations
The uncertainty constraint serves as a balancing factor between the optimization direction and the initial results. When dealing with joints that already have accurate initial estimates, two scenarios arise: (a) relying solely on the projection constraint might deteriorate the results. In this setting, our uncertainty constaint can effectively alleviate this problem; (b) relying solely on the projection constraint can also bring the optimized results closer to the ground truth. In this case, the uncertainty constraint will limit the optimization process.

### v2

@@ VI-D Ablation Study
Figure 5: Combination of a histogram representing the uncertainty for each joint and a line chart depicting the mean error (MPJPE) for each joint. The left y-axis denotes the magnitude of uncertainty, while the right y-axis signifies the error value in MPJPE.
@@
Figure 7: Visualization of the iterative optimization process. The red pose is the ground truth and the blue pose is our prediction. Color intensity reflects the number of iterations, with darker colors indicating more iterations. The initial 3D pose generated by the pre-trained model (column 3) progressively improves through iterative optimization (columns 4-7). “P” indicates the use of only the projection constraint, “P+U” denotes the combined use of projection and uncertainty constraints.
Inference speed of different Optimization Strategies. As is well known, optimization-based methods are more time-consuming. We evaluated the inference speed of various optimization strategies on a single frame using one GeForce RTX 3090, with results shown in Table IX. The baseline models include our GUMLP and the human mesh reconstruction method SPIN [18]. Although these baselines differ, their inference times are relatively fast and have minimal impact on overall runtime. Since the primary computational cost lies in the optimization stage, we can fairly compare the runtime of different methods. Our baseline model, GUMLP, reaches 310 fps. By setting the number of iterations to 5, our UAO framework reaches 48 fps, surpassing all the other optimization methods. This improvement is attributed to our UAO framework optimizing the latent state $\mathbf{z}$ rather than the entire model parameters, thereby accelerating the inference process. It is worth noting that the output of any layer within the network can be selected as the latent state for optimization. For faster processing while maintaining satisfactory performance, we select the latent state before the final fully connected layer as the optimized variable, denoted as UAO*. This choice significantly enhances the inference speed, achieving up to 125 fps.
TABLE IX: Inference speed of optimization methods. UAO* signifies employing the network’s latent state as the variable for optimization.
   Method    FPS
   GUMLP (Ours)    310.0
   SPIN [18]    60.2
   SPIN + BOA [17]    1.2
   SPIN + DynBOA [16]    0.9
   SPIN + DAPA [76]    2.3
   SPIN + CycleAdapt [22]    13.5
   SPIN + ESCAPE [21]    43.5
   GUMLP + UAO (Ours)    48.0
   GUMLP + UAO* (Ours)    125.0
VI-E Uncertainty and Error for each Joint
@@
VI-F Generalization to Unseen Camera Views
TABLE X: Performance across different views on Human3.6M in MPJPE. In-domain refers to testing under the same view as the training, while out-of-domain refers to testing under views other than the training view. $P_{i}$ denotes the $i$th camera view.
@@ VI-F Generalization to Unseen Camera Views
Table X presents the relative gains of the proposed approach over in- and out-domain cases on Human3.6M, where four camera views are available. It can be inferred that the UAO framework can not only help the in-domain cases but also bring more significant improvements to unseen camera views. This verifies its ability to break the domain gap and meet our design expectations. Besides, from the last column of the table, a large training set can bring more optimization gains, which may be explained by better pre-training leading to a better depiction of the 3D pose manifold.
VI-G Qualitative Results
@@
The visualization results on the Human3.6M dataset are shown in Figure 6. We can see that our original GUMLP is marginally superior to the previous state-of-the-art method, MGCN [7]. Moreover, employing the UAO framework brings better generation with more precise poses. For example, the 3D pose of photoing action (row 2) predicted by GUMLP with UAO (col 4) is similar to the ground truth. As a comparison, the left arms predicted by MGCN and the GUMLP are at a lower position. We further visualize the iterative process of the UAO framework under different constraints, as shown in Figure 7. The results in Figure 7 (row 1) show that relying solely on the projection constraint can cause some well-estimated joints to deviate from the ground truth. However, for the same 2D pose, incorporating the uncertainty constraint (Figure 7 (row 2)) results in an optimized pose that better aligns with the ground truth. This further validates the effectiveness of the optimization process and highlights the necessity of our uncertainty constraint.
To further demonstrate the performance in outdoor scenes, we present the qualitative results on the 3DPW test set using 2D pose detected by HRNet [47], as illustrated in Figure 8. The results indicate that the initial 3D outputs generated by GUMLP are significantly enhanced when using the UAO framework. Figure 9 also shows some occluded samples. Although our optimization strategy (UAO) is not specifically designed for occluded poses, it still performs well even in the presence of occluded body parts.
Figure 8: Visualization results on 3DPW using 2D pose detected by HRNet [47]. The red and blue poses in the Input image represent the ground truth and detected 2D poses, respectively. In the 3D pose columns (2–3), the red pose indicates the ground truth 3D pose, while the blue 3D pose corresponds to the predicted results based on the detected 2D input.
Figure 9: Occluded samples in 3DPW with 2D poses detected by HRNet [47]. The red and blue poses in the Input image represent the ground truth and detected 2D poses, respectively. In the 3D pose columns (2–3), the red pose indicates the ground truth 3D pose, while the blue 3D pose corresponds to the predicted results based on the detected 2D input.
VII Conclusion
@@
In this paper, we introduced an Uncertainty-Aware testing-time Optimization (UAO) framework for 3D human pose estimation. During training, our proposed GUMLP network estimates 3D poses along with uncertainty values for each joint. At test time, the UAO framework freezes the pre-trained network parameters and optimizes a latent state initialized by the input 2D pose. To effectively constrain the optimization in both 2D and 3D spaces, we apply projection and uncertainty constraints. Extensive experiments show that our approach achieves state-of-the-art results on popular datasets and demonstrates superior generalization in challenging outdoor scenes.